# Deciphering the global production network from cross-border firm transactions

### v2

@@ Results
Figure 4: European industrial nations and China dominate critical intermediate products, while industrial complexity is correlated with embeddedness in densely connected supply chains. [A] Various types of machinery, common components such as pumps, blades, taps and other common ’ingredients’ have high betweenness centrality. [B] European industrial nations such as Germany, Italy and Austria, as well as China, dominate high BC products. [C] Node ‘hub score’ is similar to eigencentrality in in the sense that it captures ’neighbour of neighbour’ linkages. [D] The Product Complexity Index (rank) correlates 0.46 with hub score. [E] Economic Complexity Index (ECI) has a high rank correlation of 0.72 with mean hub score. Hence, countries with high industrial complexity tend to export products which are embedded in densely connected global supply chains.
The supply chain network correlates with other efforts to reconstruct inter-product linkages. First, we consider a recent effect to construct a similar type of product network using LLM queries (Fetzer et al., 2024). In essence, the authors asked an LLM what are the top input products for each product, and constructed a network based on its answers in a similar fashion to above.
Figure 5 A shows that the (binary unweighted) AI network is much sparser than our network (see adjacency matrix in inset). Its clear from the distribution of white nodes (those with zero in-degree) that the AI network tends to have much fewer edges in sectors such as chemicals, plastics and food. Overall, our network (the SC network) has 4 times as many links as the AI network, but 5.8 times the number of links in the SC food, 4.5 times the number of links in the SC chemicals community and 3.3 times the number of links in the metals and electric machinery community. We compare communities for the AI and SC network in Figure 5 C. We find that the AI and SC product networks have similar meso-scale structure, with the exception of much larger and stronger chemicals and food clusters in the SC network.
Next, we compute a simple correlation of the full edge set of AI network and SC network, finding they are correlated up to approx 0.35 for low values of firmcount (i.e., including uncertain edges). The in- and out-degree edge correlation is significantly higher, peaking at approx 0.68 for in-degree. In general, these correlations peak for low firmcount in the range 2-4 - i.e., we keep weak edges for a dense network. Hence, while the networks look visually very different, they display non-trivial similarity in both meso structure and in- and out-degree pattern.
We also find similarity to a NAFTA derived product-level IO network. We compare our SC network to a NAFTA derived product-level ’IO network’ from Conconi et al. (2018) which was constructed based on NAFTA ’Rules of Origin’ which aims to identify the origin country of goods in order to determine if tariffs apply. In this case we find a much lower edge correlation of 0.15, a binary in-degree correlation of 0.28 and out-degree correlation of 0.32. Despite the lower correlations, however, the aggregate meso structure retains similarities to the SC (and AI) networks with food, textiles, chemicals and metals/machinery clusters (see Figure S3).
@@
Figure 5: The product network correlates with other efforts to construct inter-product links or chains. [A] The ’AI network’ of Fetzer et al. (2024) is constructed via LLM queries. Nodes with zero in-degree are shown in white. [B] The AI network is much sparser than the SC network and tends to have much fewer edges in sectors such as chemicals, plastics and food. [C] The AI and product network have similar meso-scale structure, with the exception of a much larger and stronger chemicals and food clusters in the SC network. [D-E] The correlation of the full edge set of AI network and SC network is approx 0.34, while the in and out-degree edge correlation is significantly higher, peaking at approx 0.68 for in-degree (for low firmcount in the range 2-4).
The product network picks up links identified in manually constructed single-sector supply chains. Up until now, most detailed data on inter-product linkages has come from a large number of detailed manual studies of sector specific supply chains. Here we take a selection of these, and investigate to what extent we pick up similar linkages using our method. We focus on two key sectors, electric vehicle batteries and semiconductors. For each sector, we take the set of products identified in the manual supply chains, and visualise the entries of the adjacency matrix of the subgraph of these products in Figure 1 A/C. In other words, we show the edge pattern that we find using our method for the identified product sets. We observe that we capture a high density of links between products identified in electric vehicle battery and semiconductor supply chains. While products in electric vehicle battery are highly clustered in the centre of the network, semi-conductors exhibit two distinct clusters composed of chemicals and machinery.
We wish to show that statistically its very unlikely we could have found these links at random. To do this, we simulate 100k random networks with the same number of nodes (as those identified in the sector supply chain) and compare the modularity score of our subgraph to those of the random networks. The modularity score essentially tells us how modular or connected the subgraph is compared to what would be expected at random under a null model (we use a standard null model which randomly allocates edges while keeping the node degree fixed). We compare the modularity of our network (far right axis) to the distribution of modularity scores for the ensemble of 100k subgraphs, finding an extremely low probability that our subgraph would be simulated at random (pvalue shown).
@@
Figure 6: The product network picks up links identified in manually constructed single-sector supply chains. [A/D] We capture a high density of links between products identified in electric vehicle battery and semiconductor supply chains. [B/E] While products in electric vehicle battery are highly clustered in the centre, semi-conductors exhibit two distinct clusters composed of chemicals and machinery. [C/F] To show we pick up linkages between these products that are far from random, we simulate 100k random networks with the same number of nodes and compare the modularity score (connectivity) of our subgraph (see A/D) to the random networks. We observe that our subgraphs are in the very extremes of the distributions suggesting they would be extremely unlikely to occur at random.
@@
Both downstream and upstream links predict country-product export diversification patterns. Finally, we probe whether metrics derived from our network are predictive of country-product diversification patterns. To do this, we use a similar approach to a large literature on density metrics in EEG (Hidalgo et al., 2018) to construct a ’downstream’ (and analogous ’upstream’) density metric. As illustrated in Figure 7 A/D, essentially, for a product $p$, this captures the share of top in-degree neighbouring products present in country $c$:
@@
where $J_{p}$ is the set of top $k=50$ edges from $p$ to all nodes, and $M_{p,c}=1$ if product $p$ is present in country $c$ (and otherwise 0). We construct an equivalent ’upstream density’ metric.
Similar to O’Clery et al. (2021), we perform a standard Probit regression for the probability of a product appearance of the form:
@@
where $R_{p,c}=1$ if product $p$ was present in country $c$ in 2021 but absent in 2016, $\Phi$ is a normal cumulative distribution function, and $\gamma_{p}$ and $\eta_{c}$ are product and country fixed effects respectively.
The tables in Figure B/E show that both downstream and upstream links predict country-product export diversification patterns with high predictive power (the AUC or ’area under the curve’ measures the model’s ability to distinguish between positive and negative cases) for both models with and without fixed effects. The AUC values we find are in line with the literature on predicting product entries in countries, and other similar efforts to predict diversification on a range of scales using density metrics derived from a variety of networks capturing ’capability overlap’ between products (Hidalgo et al., 2007; O’Clery et al., 2021).
Figure C/F shows, perhaps counter-intuitively but in line with some of the validation steps above, that the predictive power is highest for low firmcount. Hence, including ’weak links’ for which there is relatively lower certainty is informative for the model. The predictive power is also higher for lower entry threshold - i.e., we better predict small increases in Rpop (which compares a country’s per capita production levels in a product to the world’s overall per capita production of the product) over a 5 year period compared to larger jumps.
Figure 7: Both downstream and upstream links predict country-product export diversification patterns. [A/D] Downstream/upstream-density captures a country’s presence in products that are downstream/upstream of the target product. [B/E] Both downstream and upstream links predict country-product export diversification patterns with fairly high predictive power (AUC) for cases with and without fixed effects. [C/F] The predictive power is highest for low firmcount (i.e., including ’weak links’), and for low rpop entry threshold (i.e., we predict small increases in rpop (product concentration relative to GDPPC).
@@ Discussion
In this paper we develop a highly granular network tool to describe complex global inter-related product supply chains. We explore its broader structural characteristics, and investigate which countries dominate common input nodes with high betweenness, and which are active in highly embedded ’hub’ nodes. Despite an absence of ’ground truth’ data, we validate it in several ways including comparing it to a variety of other efforts to construct product level supply chain networks, and manually constructed data on sector specific supply chains.
@@
There are a wide range of potential research and policy uses for this type of tool. From a policy viewpoint, this tool could be used to, for example, quickly map supply chains for new products which have not already been mapped manually, as well as capture inputs not typically considered in manual mapping such as machinery and tool inputs. More broadly, when augmented with country level information, the tool can be used to gain insights into the processes underlying globalisation and the ’unbundling’ of supply chains (Baldwin, 2006).
@@
The network structure provides key insights on ’central nodes’ and clusters of highly interconnected products. Central nodes might be worthy of particular focus or protection. For example, the fact that steel has very high betweenness centrality indicates that it likely plays a role in many paths (chains) and might be seen as a critical good. This accords with the fact that the UK government moved aggressively to protect domestic production of domestic virgin steel in early 2025 in the face of impending closure of British Steel. More broadly, the network can be used to quantify supply chain vulnerabilities and model the propagation of shocks and policies such as tariffs in a more detailed, comprehensive manner than previously possible. In the case of product clusters, we can see network communities as a sort of alternative product hierarchy relative to official sectoral classifications. Highly connected clusters would naturally be collectively impacted by many kinds of cascading shocks and so they might be collectively targetted by policy interventions rather than individually.
@@
In terms of wider applications, there is a growing literature on ’embodied carbon’ as carbon from input products is incorporated to final goods (Sato, 2014). Little is known, however, about so-called Scope 3 emissions - or those transmitted via supply chains - despite the fact they thought to be growing faster than other types of emissions (Hertwich and Wood, 2018). This type of detailed network which captures - at least in theory - the full set of product relations could be used to track ’product life-cycle’ emissions as products are manufactured and then used in other manufacturing processes (Wang and Gupta, 2011).
@@
Here we show that country-product export diversification is associated with both forward and backward linkages, with a slightly stronger signal for backward linkages. The network could be used as a policy tool to identify opportunities in the form of upstream and downstream products which could be targets for industrial policy support. This type of policy approach is used, for example, in ’Smart Specialisation’ efforts at a European level in which regions aim to specialise and diversify in related sectors (Balland et al., 2019). Beyond nodes, it also sheds light on potential new edges. For example, it could be used to identify linkages which exist globally but not locally.
@@
There are some limitations to our approach, as well as potential sources of noise. In order to build our network, we aggregate data over 3 years based on advice from the supplier and our analysis of the raw firm-interaction data that suggests the data is most complete for these latter years. This approach, however, does not currently allow for temporal analysis. Future work may investigate the extent to which disaggregating into distinct time periods and recomputing the network for each is feasible. Secondly, since our transactions are between entities located in different countries, we do not capture local linkages. For example, if the water used to produce beer is predominantly purchased locally, we may not capture it as a key input. Hence, we over-emphasise inputs and relationships in global value chains rather than local ones.
@@
Third, there is no doubt some noise in our data due to the uneven composition of the firm-level data. For example, while we have good coverage of firms across developed countries, we by no means have the ’universe’ of firms in our dataset, and some countries/sectors will be more represented than others. This issue is less problematic for our tool than say country-specific or firm-level analysis, but nonetheless it will introduce noise which we cannot fully mitigate against. Fourth, there may be other sources of noise, such as short term changes in purchasing patterns. For example, we noticed an unusual input to semiconductors - washing machines. But it turned out that due to a huge shortage in input components in 2022/23, semiconductor firms were actually buying washing machines to strip them of their chips. Another source of noise is likely to be multi-product firms. While we have taken efforts to reduce this, by omitting firms that sell products in more than five HS Sections, and looking for common purchases by all firms that make a product, there will no doubt be some noise remaining due to this issue.
@@
Fifth, there may be idiosyncratic differences between supply chain relationships within regions that we aggregate over by computing a network using all firm transactions. For example, battery supply chains may look different in the US vs Asia. While we have looked at some geographic variations here (see SI), in future work we may reconstruct the network using data for specific regions to further investigate this issue. Finally, we present the network at HS 4-digit level. It may be possible in future work to construct the network at 6-digit level, but due to an exponential increase in the number of edges to estimate, we have much fewer number of firms (firmcount parameter) for each linkage and a more complicated validation exercise with fewer sources of external data to validate against. Again, we leave this to future work.
Methods
@@
The cut of data we used from GSCIP on firm-level transactions covers the time period between Jan 2021 and Dec 2023.
We take the following processing steps:
@@
  • We aggregate all firms to an ’owner-country’ level. In other words, we have Coca Cola-Japan and Coca Cola-Belgium.
Construction of supply chain network
A concern about our approach might be that it could pick up spurious links due to the presence of firms in our dataset which make many products (and hence we cannot infer specific inputs to products from their purchases). Our methodology is designed to reduce this issue since we look at the purchases of not just one firm but all firms that make a specific product. To mitigate this potential issue further, we drop all firms that export products in five or more HS Sections (e.g., this would exclude wholesalers and multi-nationals with large portfolios of products such as Proctor and Gamble).
@@
Community detection We use the Stability algorithm of Delvenne et al. (2010) to detect communities in the SC and AI networks. The algorithm detects partitions on a range of scales corresponding to random walker time on the network. The higher the resolution, the longer the time, the larger (and fewer) the communities. We use the ’full’ (non-linearised) specification for a directed Laplacian, with 100 iterations. In the SI we show the number of communities, variation of information across the iterations (a low value tells us we find similar partitions for different random seeds), and the stability score for our set of partitions corresponding to a range of timescales.
Modularity score We use the Modularity score to compare the connectivity of the subgraph induced by each of our sector-specific supply chains (EV batteries and semi-conductors) to 100k randomly generated subgraphs of the same size. Modularity (Newman, 2006) compares the connectivity of a community to what would be expected under a random null model - in this case a network in which the node degree is preserved but the edges randomised.
Effectively, we decompose the directed Modularity score proposed by Leicht and Newman (2008) to evaluate a set of communities for a full network, and use it instead to evaluate the connectivity of a subgraph (i.e., a single community). This is mathematically coherent as Modularity is a sum over the connectivity within communities. Specifically, we compute:
@@
To capture product entries, we wish to compute a measure of a countries’ concentration in a product. The traditional approach is Revealed Comparative Advantage (RCA) proposed by Balassa (1965) captures is the proportion of the country’s exports in a product divided by the global proportion of exports in a product. We use a related approach, denoted Rpop, which helps remove noise due to fluctuations in commodity prices, seasonal employment or currency exchange rates (Hidalgo, 2021). Specifically, it compares a country’s per capita production levels in a product to the world’s overall per capita production of the product:
@@
We conduct the econometric analysis for a set of 235 countries. We drop products not in the main connected component of the network and those with relatively small global trade (less than 2 billion\$ in 2016) for a total of 683 products. Of a possible 235x683 observations, 86% were absent ($Rpop<0.05$) in 2016. Of these, 3.5% of products appeared by 2021 (with $Rpop>0.1$). While we keep the absence threshold fixed, we show results for a range of presence thresholds in Figure 7 C and F.
Supplementary Information
Overview of underlying data

### v3

@@ Results
Figure 4: European industrial nations and China dominate critical intermediate products, while country-level industrial complexity is correlated with embeddedness in densely connected supply chains. [A] Various types of machinery, common components such as pumps, blades, taps and other common ’ingredients’ are ’chokepoints’ with high betweenness centrality (BC). [B] European industrial nations such as Germany, Italy and Austria, as well as China, dominate high BC products. [C] Node ‘hub score’ is similar to eigencentrality in in the sense that it captures dense ’neighbour of neighbour’ linkages. [D] The Product Complexity Index (rank) correlates 0.46 with hub score. [E] Economic Complexity Index (ECI) has a high rank correlation of 0.85 with mean hub score. Hence, countries with high industrial complexity tend to export products which are embedded in densely connected global supply chains.
European industrial nations and China dominate critical intermediate products, while industrial complexity is correlated with embeddedness in densely connected supply chains. Moving to node-level metrics, shown in SI Figure S2, we find that machinery products, including those related to cars and motors, tend to have a large in-degree (they require many inputs), while components, plastics and packaging tend to have high out-degree (i.e., they are inputs to many products). Further, consistent with what we might expect, deploying the UN BEC classification that sorts products into capital, intermediate and consumption goods, we find that capital products are inputs to many products (high out-degree), while both capital and consumption goods have high in-degree – they require many ingredients (SI Figure S2).
From a network structure perspective, betweenness centrality (BC) measures the extent to which a node acts as a key ’juncture’ through which many paths (i.e., supply chains) cross. Mathematically, the betweenness centrality of a node is the number of times a path (between any pair of nodes) transits that node. In Figure 4 A, we highlight products with high betweenness centrality, including various types of machinery, common components such as pumps, blades, taps and motors, plastics, packaging, and chemical binders. Nations which produce these key inputs have an important role in global supply chains, controlling so-called ’choke points’. To investigate this, we compute the Revealed Comparative Advantage (a measure of export concentration of a country in a product) for each country-product pair from global trade data (Comtrade 2024). In Figure 4 B, we show a global map with countries shaded by their share of top-50 BC products exported with RCA. We find European industrial nations such as Germany, Italy and Austria, as well as China, dominate high BC products.
Next we investigate the extent to which highly industrialised nations engage export goods which are highly embedded in densely inter-connected global supply chains. To capture production complexity, we adopt an approach proposed by [36] to infer product complexity (PCI) from patterns of geographical production captured in trade data. To measure the localised connectivity and embeddedness of product in global supply chains, we compute the hub score of each node [41]. This is similar to eigencentrality in the sense that it captures ’neighbour of neighbour’ linkages, but is suitable for directed networks such as ours. We find that the PCI (obtained for 2023 from the Harvard Atlas of Economic Complexity) has a rank correlation of 0.46 with node hub score. In essence, complex products are highly embedded in global supply chains. At the country level, the Economic Complexity Index (ECI), which is mathematically equivalent to the mean PCI of products exported with RCA, has a high rank correlation of 0.85 with mean hub score. Hence, countries with high industrial complexity tend to export products which are highly embedded in densely inter-connected global supply chains.
@@
Figure 5: The product network correlates with other efforts to construct inter-product links or chains. [A] The ’AI network’ of [37] is constructed via LLM queries. Nodes with zero in-degree are shown in white. [B] The AI network is much sparser than the SC network and tends to have much fewer edges in sectors such as chemicals, plastics and food. [C] The AI and product network have similar meso-scale structure, with the exception of a much larger and stronger chemicals and food clusters in our network (SC network). [D-E] The correlation of the full edge set of AI network and SC network is approx 0.35, while the in and out-degree edge correlation is significantly higher, peaking at approx 0.68 for in-degree (for low firmcount in the range 2-4).
The supply chain network correlates with other efforts to reconstruct inter-product supply chain linkages. While there are no ’ground truth’ datasets for us to validate against, and concordances between detailed products and aggregate industries are highly noisy limiting utility of comparing to industry-based Input-Output-type datasets, here we compare our network to a recent attempt to construct a similar type of product network using LLM queries [37]. Specifically, the authors asked an LLM what are the top input products for each product, and constructed a (binary unweighted) network based on its answers.
Figure 5 A and B show that the AI network is in general much sparser than our network (we keep the same node layout as above). Its clear from the distribution of white nodes (those with zero in-degree) that the AI network tends to have much fewer edges in sectors such as chemicals, plastics and food. Overall, our network (the ’SC’ network, with parameters set as above) has 4 times as many links as the AI network, but 5.8 times the number of links in the SC food community, 4.5 times the number of links in the SC chemicals community and 3.3 times the number of links in the metals and electric machinery community. We compare communities for the AI and SC network in Figure 5 C. We find that the AI and SC product networks have a similar meso-scale structure, with the exception of much larger and more strongly connected chemicals and food clusters in the SC network.
Next, we compute a simple correlation of the full edge set of AI network and SC network, finding they are correlated up to approx 0.35 for low values of firmcount (i.e., including uncertain edges). The in- and out-degree edge correlation is significantly higher, peaking at approx 0.68 for in-degree. In general, these correlations peak for low firmcount values in the range 2-4 - i.e., we keep weak edges for a dense network. Hence, while the networks look visually very different, they display non-trivial similarity in both meso structure and in- and out-degree pattern.
We also compare our SC network to a product network constructed based on NAFTA ’Rules of Origin’ [38] which aims to identify the origin country of goods in order to determine if tariffs apply. In this case we find a much lower edge correlation of 0.15, a binary in-degree correlation of 0.28 and out-degree correlation of 0.32. Despite the lower correlations, however, the aggregate meso structure retains similarities to the SC (and AI) networks with food, textiles, chemicals and metals/machinery clusters (see SI Figure S4).
@@
Figure 6: The product network has dense connections between products identified in manually constructed single-sector supply chains. [A/D] We capture a high density of links between products identified in electric vehicle battery and semiconductor supply chains. [B/E] While products in electric vehicle battery are highly clustered in the centre, semi-conductors exhibit two distinct clusters composed of chemicals and machinery. [C/F] To show that we pick up a pattern of dense linkages between these products which is far from random, we simulate 100k random networks with the same number of nodes and compare the modularity score (connectivity) of our subgraph (see A/D) to the random networks. We observe that our subgraphs are in the very extremes of the distributions suggesting they would be extremely unlikely to occur at random.
The product network has dense connections between products identified in manually constructed single-sector supply chains. Up until now, most detailed data on inter-product linkages has come from a large number of detailed manual studies of sector-specific supply chains. Here we take a selection of these, and investigate to what extent we pick up similar patterns using our method. We focus on two key sectors, electric vehicle batteries [10] and semiconductors [9]. For each sector, we take the set of products identified in the manual supply chains, and visualise the entries of the adjacency matrix of the subgraph of these products in Figure 6 A/D. In other words, we show the edge pattern that we find in our network for the manually identified product sets. We observe that we capture a high density of links between products identified in electric vehicle battery and semiconductor supply chains. While products in electric vehicle battery are highly clustered in the centre of the network, semi-conductors exhibit two distinct clusters composed of chemicals and machinery products.
Could these links be random? In other words, if we selected a random set of products, would they also have a similar link density? To test this, we simulate 100k random subgraphs with the same number of nodes (as those identified in the sector supply chain) and compare the modularity score of our subgraph to those of the random subgraphs. The (directed) modularity score [42] essentially tells us how modular or connected the subgraph is compared to what would be expected at random under a null model (we use a standard configuration null model which randomly allocates edges while keeping the node degree fixed). We compare the modularity of our network (far right axis) to the distribution of modularity scores for the ensemble of 100k random subgraphs, finding an extremely low probability that our subgraph would be simulated at random (p-value shown).
Figure 7: Upstream links predict country-product export diversification patterns for countries with higher industrial complexity. [A/E] Both downstream- and upstream-density captures a country’s presence in products that are downstream/upstream of the target product. [B-D] The downstream model (no FE) performs well, in particular for low firmcount (many edges) and top 100 ECI countries. [F-H] The upstream model (no FE) performs consistently very well over a wide parameter space, with highest values for low firmcount (many edges), low increase in Rpop, and top 100 ECI countries. [I/J] We show the full regression table (include FE) for both downstream and upstream models for high ECI countries and parameters firmcount equals 1 and entry Rpop threshold 0.1.
@@
Upstream links strongly predict country-product export diversification patterns for countries with higher industrial complexity. Finally, we probe whether metrics derived from our network are predictive of country-product diversification patterns. To do this, we use a similar approach to a large literature on density and ’relatedness’ metrics in the Economic Complexity and Evolutionary Economic Geography fields [43]. Specifically, we construct a ’downstream’ (and analogous ’upstream’) density metric, as illustrated in Figure 7 A/E, for a product $p$ in country $c$ which captures the share of top in-degree neighbouring products of $p$ present in country $c$:
@@
where $J_{p}$ is the set of top $k=40$ edges from $p$ to all nodes, and $M_{p,c}=1$ if product $p$ is present in country $c$ (and otherwise 0). We also construct an equivalent ’upstream density’ metric which considers upstream neighbouring nodes.
Similar to [35], we perform a standard Probit regression for the probability of a product appearance of the form:
@@
where $R_{p,c}=1$ if product $p$ was present in country $c$ in 2021 but absent in 2016, $\Phi$ is a normal cumulative distribution function, and $\gamma_{p}$ and $\eta_{c}$ are product and country fixed effects respectively. Rpop, which compares a country’s per capita production levels in a product to the world’s overall per capita production of the product, is a threshold used to define product ’presence’ or ’absence’ (similar to Revealed Comparative Advantage [44]).
In Figure 7 B-D and F-H, we show the predictive power (the AUC or ’area under the curve’) for both models (downsteam and upstream respectively) without fixed effects for three distinct country sets: a full set of 235 countries, the top 100 countries by ECI and the bottom 100 countries by ECI. The grids vary parameters firmcount and Rpop. Figure 7 J and I show the full econometric model (with various fixed effect models) in table format for the subset of 100 high ECI countries with network parameters firmcount equals 1 and entry Rpop threshold 0.1.
Overall, across variation in model parameters and country subsets, we find that both downstream and upstream linkages are predictive of diversification patterns - but that upstream linkages are significantly more predictive especially for the higher ECI group of countries. This is consistent with an established literature on the role of upstream linkages in diversification processes [23, 32, 33]. The high ECI set is overall better predicted relative to the full set and low ECI set, which is likely due to the fact that, as we saw above, these countries are more strongly embedded in densely connected global value chains.
It is also clear that, perhaps counter-intuitively, the predictive power is highest for low firmcount values. Hence, including ’weak links’ for which there is relatively lower certainty is informative for the model. We are also better in general at predicting small increases in Rpop over a 5 year period compared to larger jumps in Rpop. However, many model specifications do achieve high AUC for larger values of Rpop, e.g., for Rpop=0.5, the upstream model (for high ECI countries/low firmcount) has AUC close to 0.8.
The AUC values we find are in line with (or exceed) the literature on predicting product entries in countries, and other similar efforts to predict diversification using density metrics derived from a variety of networks capturing ’capability overlap’ or linkages between products [45, 35, 46]. In the SI we show similar results for a 10 year model (2011-21), and variation in density parameter $k$ (this peaks around $k=40$ for most specifications, which is consistent with previous work [35, 46], and used in Figure 7 here).
@@ Discussion
In this paper we develop a highly granular network tool to describe complex global product supply chains. We explore its broader structural characteristics, and investigate which countries dominate chokepoint nodes with high betweenness, and which are embedded in highly connected supply chains. Despite an absence of ’ground truth’ data, we compare the network to a variety of other efforts to construct product-level supply chain networks, as well as manually constructed sector-specific supply chains.
@@
There are a wide range of potential research and policy uses for this type of tool. From a policy viewpoint, this tool could be used to, for example, gain an overview of the location of individual products in global supply chains in the absence of manual mapping, as well as capture inputs not typically considered in manual mapping such as machinery and tool inputs. More broadly, when augmented with country level information, the tool can be used to gain insights into the processes underlying globalisation and the ’unbundling’ of supply chains [47]. These dynamics have implications for economic inter-dependence between countries, and the increase the potential for economic tools like tariffs and subsidies to exert pressure and influence [48].
@@
The network structure provides key insights on ’central nodes’ and clusters of highly interconnected products. Central nodes might be worthy of particular focus or protection. For example, the fact that steel has very high betweenness centrality indicates that it likely plays a role in many paths (chains) and might be seen as a critical good. This accords with the fact that the US government moved aggressively to protect domestic production of steel throughout 2025 via steep tariffs on imports of both steel and steel ’derivatives’ which includes products with a high steel content such as bolts and bicycles. More broadly, the network can be used to quantify supply chain vulnerabilities and model the propagation of shocks - ranging from financial shocks to conflict to natural disaster - and the impact of policies such as tariffs in a more detailed, comprehensive manner than previously possible. For example, highly connected clusters would naturally be collectively impacted by many kinds of cascading shocks and so they might be jointly targetted by policy interventions rather than individually.
@@
In terms of wider applications, there is a growing literature on ’embodied carbon’ as carbon from input products is incorporated to final goods [49]. Little is known, however, about so-called Scope 3 emissions - or those transmitted via supply chains - despite the fact they thought to be growing faster than other types of emissions [50]. This type of detailed network which captures global production paths could be used to track cross-border ’product life-cycle’ emissions as products are manufactured and then exported to be used in other manufacturing processes [51].
@@
Here we show that country-product export diversification is associated with both forward and backward linkages, with a stronger signal for forward (upstream) linkages, and more complex economies. Consistent with the view that upstream products, which are inputs to many other goods, are often highly complex and high value-add, such as semi-conductors or optical microscopes, it seems that recent diversification patterns tend to move in this direction. This is consistent with previous findings [32]. In a similar vein to [45, 22], the network could be used as a policy tool to identify ’feasible’ opportunities in the form of both upstream and downstream products which could be targets for industrial policy support. This type of policy approach is used, for example, in ’Smart Specialisation’ efforts at a European level in which regions aim to specialise and diversify in related sectors [52].
@@
There are some limitations to our approach, as well as potential sources of noise. Firstly, in order to build our network, we aggregate data over 3 years based on advice from the supplier and our analysis of the raw firm-interaction data that suggests the data is most complete for these latter years. This approach, however, does not allow for temporal analysis. Future work may investigate the extent to which disaggregating into distinct time periods and recomputing the network for each is feasible. Secondly, since our transactions are between entities located in different countries, we do not capture local linkages. For example, if the water used to produce beer is predominantly purchased locally, we may not capture it as a key input. Hence, we over-emphasise inputs and relationships in global value chains rather than local ones.
@@
Third, there is no doubt some noise in our data due to the uneven composition of the firm-level data. For example, while we have good coverage of firms across developed countries, we by no means have the ’universe’ of all global firms in our dataset, and some countries/sectors will be more represented than others. This issue is less problematic for our tool than say country-specific or firm-level analysis, but nonetheless it will introduce noise which we cannot fully mitigate against. Fourth, there may be other sources of noise, such as short term changes in purchasing patterns. For example, we noticed an unusual input to semiconductors - washing machines. It turned out that due to a huge shortage in input components in 2022/23, semiconductor firms were actually buying washing machines to strip them of their chips [53]. Another source of noise is likely to be multi-product firms. While we have taken efforts to reduce this by omitting firms that sell products in more than five HS Sections, and our overall approach based on common purchases by all firms that make a product, there will no doubt be some noise remaining due to this issue.
@@
Fifth, there may be idiosyncratic differences between supply chain relationships within regions that we aggregate over by computing a network using all firm transactions. For example, battery supply chains may look different in the US vs Asia. In future work we may reconstruct the network using data for specific regions to further investigate this issue. Finally, we present the network at HS 4-digit level. It may be possible in future work to construct the network at 6-digit level, but due to an exponential increase in the number of edges to estimate, we have much fewer number of firms (firmcount parameter) for each linkage and fewer sources of external data to compare the network structure to. Again, we leave this to future work.
Materials and Methods
@@
Our GSCIP data includes firm-level transactions during the time period between January 2021 and December 2023. We take the following processing steps:
@@
  • We aggregate all firms to an ’owner-country’ level (see SI). In other words, we have Coca Cola-Japan and Coca Cola-Belgium.
Multi-product firms
A concern about our approach might be that it could pick up spurious links due to the presence of firms in our dataset which make many products (and hence we cannot infer specific inputs to products from their purchases). Our methodology is designed to reduce this issue since we look at the purchases of not just one firm but all firms that make a specific product. To mitigate this potential issue further, we drop all firms that export products in five or more HS Sections (e.g., this would exclude wholesalers and multi-nationals with large portfolios of diverse products such as Proctor and Gamble).
@@
Community detection We use the Stability algorithm of [39] to detect communities in the SC and AI networks. The algorithm detects partitions on a range of ’scales’ corresponding to the ’time’ spent by a random walker exploring the network. The lower the time, the smaller (but more numerous) the communities detected. The longer the time, the larger (and fewer) the communities. We use the ’full’ (non-linearised) specification for a directed Laplacian, with 100 iterations. In the SI we show the number of communities, variation of information across the iterations (a low value tells us we find similar partitions for different random seeds), and the stability score for our set of partitions corresponding to a range of timescales.
Modularity score We use the Modularity score to compare the connectivity of the subgraph induced by each of our sector-specific supply chains (EV batteries and semi-conductors) to 100k randomly generated subgraphs of the same size. Modularity [54] compares the connectivity of a community to what would be expected under a random null model - in this case a network in which the node degree is preserved but the edges randomised.
Effectively, we decompose the directed Modularity score proposed by [42], which aims to evaluate the connectivity of a set of communities in a network, and use it instead to evaluate the connectivity of a single subgraph (i.e., a single community). This is mathematically coherent as Modularity is a sum over the connectivity within communities. Specifically, we compute:
@@
To capture product entries, we wish to compute a measure of a countries’ concentration in a product. The traditional approach is Revealed Comparative Advantage (RCA), proposed by [44], captures the proportion of the country’s exports in a product divided by the global proportion of exports in a product. We use a related approach, denoted Rpop, which helps remove noise due to fluctuations in commodity prices, seasonal employment or currency exchange rates [55]. Specifically, it compares a country’s per capita production levels in a product to the world’s overall per capita production of the product:
@@
We conduct the full econometric analysis for a set of 235 countries. We drop products not in the main connected component of the network, which is dependent on the values of parameters firmcount and edge threshold, and 28 products with global trade in 2021 less than \$10m. For the tables shown in Figure 7 I/J, with both firmcount and edge threshold equal to 1, we have $n=1198$. Of a possible 100x1198 observations in Figure 7 I (high ECI country set), 76% were absent ($Rpop<0.05$) in 2016. Of these, 3.6% of products appeared by 2021 (with $Rpop>0.1$).
Acknowledgements
We would like to thank Christian Diem, Tobias Reisch, Adnan Khan and Fergus Cumming for their useful input and comments, and Shaun Hoang for his research assistance. The data and computing infrastructure was provided by the the UK Government Global Supply Chain Intelligence Programme (GSCIP).
Funding: NOC would like to acknowledge funding from the 2024-2025 UKRI Policy Fellowship programme (ES/Y004817/1_583688).
Author contributions:
Conceptualization: NOC, DTH
Methodology: NOC, BRB, TS, DTH
Investigation: NOC, BRB, TS
Visualization: NOC
Supervision: NOC, BRB, DTH
Writing-original draft: NOC
Writing-review & editing: NOC, BRB, TS, DTH
Competing interests: Any views expressed are solely those of the author(s) and so cannot be taken to represent those of the Foreign Commonwealth and Development Office (FCDO), or to state FCDO Policy. This paper should therefore not be reported as representing the views of the FCDO, its Ministers or His Majesty’s Government.
Data and materials availability: The underlying firm-level data was made available to the research team via the UK Government Global Supply Chain Intelligence Programme (GSCIP). This sensitive commercial data is held under license is, unfortunately, not available to external researchers.
Supplementary Materials for
Deciphering the global production network from cross-border firm transactions
Neave O’Clery* et al.
*Corresponding author. Email: n.oclery@ucl.ac.uk
This PDF file includes:
Supplementary Text
Figs. S1 to S6
Supplementary Text
Data